# Exploring the generalizability of the optimal 0.234 acceptance rate in random-walk Metropolis and parallel tempering algorithms

### v1

@@ 3.4.4 A Tale of Three Mixtures: Inhomogenous Scaling Factors
Table 7: Inhomogenously scaled three mixture target distribution experimental optimal RWM acceptance rates for different dimensions.
Dimension $d$ 20 30 50
Optimal Acceptance Rate 0.1740 0.2107 0.2290
[…]
Overall, the three-mixture examples and their results yield promising avenues for extension of the theory, going beyond the i.i.d. product assumption to a target density that is a weighted sum of individual Gaussian densities. Furthermore, the 0.234 figure seems to be relevant even for inhomogeneous scaling in each Gaussian in a Gaussian mixture.
@@ 4.1 Parallel Tempering Method
The parallel tempering method [8], also known as the Metropolis-coupled Markov chain Monte Carlo algorithm or the replica exchange method, is a version of the Metropolis algorithm that specialises in sampling from multimodal distributions. In this case, we sample from a target distribution $\pi(\cdot)$ with density $f_{d}(\mathbf{x})$ on a state space $\mathcal{X}$. We define a sequence of tempered target densities $f_{d}^{\beta_{j}}(\mathbf{x})$ where $10\leq\beta_{n}<\beta_{n-1}<\dots<\beta_{1}<\beta_{0}=$ 1 are the inverse temperature values. We require that $f_{d}^{\beta_{0}}(\mathbf{x})=f_{d}(\mathbf{x})$ (that is, that $\beta_{0}=1$). Typically, the tempered target densities are simply powers of the original density, and we assume this for all parallel tempering experiments. Then, we run a Markov chain for each of the $n+1$ values of $\beta$, each with its own tempered target density. Eventually, the $\beta_{0}=1$ chain, also called the “cold” chain, with its stationary density $f_{d}^{\beta_{0}}(\mathbf{x}_{0})$ should correspond to the original target density of interest $f_{d}(\mathbf{x})$, and the idea is that using information from other chains with higher temperatures (lower $\beta$) will speed up the mixing in the cold chain.
In each iteration of the algorithm, we alternate between within temperature and temperature swap moves. The former indicates that each chain takes a standard RWM update within its respective $f_{d}^{\beta_{j}}$. The latter indicates that for two chains with $\beta_{j},\beta_{k}$ inverse temperature values that are adjacent in the sequence of $\beta$ values, switch the chain values of $\mathbf{x}_{j}$ and $\mathbf{x}_{k}$ with probability
$\min\left\{1,\dfrac{f_{d}^{\beta_{j}}(\mathbf{x}_{k})f_{d}^{\beta_{k}}(\mathbf{x}_{j})}{f_{d}^{\beta_{j}}(\mathbf{x}_{j})f_{d}^{\beta_{k}}(\mathbf{x}_{k})}\right\}$ (3)
@@ 4.1.1 Optimal Scaling for Parallel Tempering
Since the hot chains can explore the state space more quickly, we would like to maximise how frequently we can swap values from the hottest chain to the coldest chain so that the cold chain can mix faster and escape local modes. To do this, we maximise the effective speed with which the chain values move along in the inverse temperature domain. The spacings of the inverse temperatures $\beta_{i}$’s are crucial to this efficiency. If $\beta_{j}$ and $\beta_{k}$ are too far apart, we usually reject these swaps described by (3), but if they are too near, the swaps will not improve mixing. We would like to swap just the right amount to maximise mixing in the temperature domain.
The expected squared jumping distance (ESJD) for parallel tempering thus refers to the expected squared jump in inverse temperatures. Formalising this, when we attempt to swap the chain values between the inverse temperatures $\beta$ and $\gamma:=\beta+\epsilon$ where $\beta,\epsilon>0$ and $\beta,\gamma\leq 1$, the swap is either accepted, in which case the values move a squared distance of $(\gamma-\beta)^{2}=\epsilon$, or the swap is rejected, in which case the distance moved is 0. This leads to a very natural definition for ESJD, which is
@@
Under the restrictions of an i.i.d. product target density $f_{d}(\mathbf{x})=\prod_{i=1}^{d}f(x_{i})$, where the tempered distributions are powers of the original density, and the inverse-temperature spread $\epsilon=d^{-1/2}\ell$ for some positive constant $\ell$, choosing $\ell$ such that the inverse temperature spacing yields a swap acceptance probability of approximately 0.234 is also optimal [1] in the sense that it maximises the ESJD of a parallel tempering algorithm.
@@ 5.1 Methodology
We conduct experiments that examine the trend of ESJD with swap acceptance rate for each of the “rough carpet” distribution from Section 3.4.1, three-mixture distribution from Section 3.4.3, and standard multivariate Gaussian distribution as a baseline for comparison. For each distribution, we run many simulations where each simulation is a parallel tempering algorithm with 40 different average swap acceptance rate values. We explain how we set up these algorithms to match these swap rates in Section 5.1.1. Each parallel tempering algorithm runs for 20,000 iterations, so every individual chain in the algorithm takes 20,000 steps. We attempt a swap every 20 steps. Furthermore, we ran each algorithm instance over 5 different seeds and average the results to reduce the effects of randomness caused by a particular seed. Overall, each experiment on a single distribution has 200 simulations which runs for 20,000 iterations per chain, and the number of chains for each parallel tempering algorithm simulation is determined by the inverse temperature spacing given by the intended average swap acceptance rate. We refer the reader to Section 5.1.1 for details on how these inverse temperatures were set up.
For each target distribution, we set the tempered target distribution to simply be the original distribution raised to the power of $\beta$. We set the modes to be far away enough such that the other modes were unreachable by the standard random-walk Metropolis algorithm after 100,000 iterations, but still reasonably close so that the parallel tempering method would show the cold chain values would swap between modes somewhat frequently.
@@ 5.1.1 Constructing an Inverse Temperature Ladder Iteratively
The most common method to construct an inverse temperature ladder selects the inverse temperatures using a geometric series spacing. However, since we are examining the 0.234 swap acceptance rule, we need to have a way of constructing the inverse temperature spacings such that the probability of a swap between adjacent chains is approximately 0.234 (or any other value). To construct an inverse temperature ladder with a desired swap acceptance rate $s$, we use an iterative procedure adapted from [1, Section 2.2]. Starting with $\beta_{0}=1$ and a minimum value $\beta_{min}=0.01$, we iteratively add $\beta$’s to the ladder. Let $\beta_{curr}$ denote the most recent $\beta$ in the ladder. We initialize $\rho_{n}=0.5$ ($n=1$ initially) and set $\beta^{*}=\beta_{curr}(1+e^{\rho_{n}})^{-1}$. To determine if $\beta^{*}$ should be added to the ladder, we draw $N$ samples (our implementation uses $N=3000$) from the target distribution tempered by $\beta^{*}$ and $\beta_{curr}$, and calculate the average swap probability $a$. If $a$ is within $s\pm 0.005$, we add $\beta^{*}$ to the ladder and set $\beta_{curr}=\beta^{*}$. Otherwise, we update $\rho_{n+1}=\rho_{n}+n^{-0.25}(a-s)$ and recalculate $\beta^{*}$. This process continues until $\beta^{*}\leq\beta_{min}$, at which point we add $\beta_{min}$ to the ladder and terminate.
5.2 Multivariate Gaussian Target Distribution

### v3

@@ 3.4.4 Simulations and Results
Table 4: Empirical optimal RWM acceptance rates for three different $d$-dimensional Rosenbrock target distributions.
Target $d=4\,\text{or}\,5$ $d=9\,\text{or}\,10$ $d=19\,\text{or}\,20$ $d=29\,\text{or}\,30$
$\pi_{8}$: Full Rosenbrock 0.2047 0.2359 0.2384 0.2337
$\pi_{9}$: Even Rosenbrock 0.0646 0.1430 0.2003 0.2107
$\pi_{10}$: Hybrid Rosenbrock 0.1394 0.2272 0.2349 0.2340
Our experiments with the Rosenbrock family of targets, presented in Table 4, demonstrate that the 0.234 heuristic remains effective even for distributions with non-linear dependencies and curvature. For the Full Rosenbrock ($\pi_{8}$) and Hybrid Rosenbrock ($\pi_{10}$) targets, the optimal acceptance rate robustly converges to approximately 0.234 as the dimension increases. Although these targets are not i.i.d. and possess a highly anisotropic banana-shaped geometry, they should satisfy the general criteria from Sherlock [38] (stated as Assumption 2.2) in the high-dimensional limit. As $d$ grows, the complex local correlations become less influential relative to the global structure, allowing the target’s geometric properties to stabilize in the manner required by Assumption 2.2. On the other hand, the Even Rosenbrock ($\pi_{9}$) has a significantly lower optimal acceptance rate that increases slowly with dimension. This seems reasonable, as $\pi_{9}$ is a product of independent 2-dimensional Rosenbrock blocks. The proposal scale can then be fit to the local geometry of each 2-dimensional block where the high-dimensional scaling theory is less applicable.
Figure 5: ESJD vs. acceptance rate for the 2-dimensional Rosenbrock target distribution $\pi_{8}$ under RWM with a Gaussian proposal. Red dotted line indicates an acceptance rate of 0.234.
The result for the 2-dimensional Rosenbrock target, shown in Figure 5, would add support for this hypothesis. Unlike the higher-dimensional cases, the ESJD does not exhibit a clear peak around a non-zero acceptance rate. Instead, it appears to increase monotonically as the acceptance rate approaches zero. This behaviour indicates that for this low-dimensional, highly anisotropic target, the highest efficiency is achieved by proposing extremely large jumps. Such jumps are almost always rejected, but the rare accepted jump is so large that it dominates the ESJD calculation.
3.5 Non-smooth Target Densities: Continuous Hypercube
Now, consider a RWM using a Gaussian proposal and a target density still of the product form $\pi(\mathbf{x})=\prod_{i=1}^{d}f(x_{i})$, but where the individual component density $f$ is Uniform$[a,b]$.
$\pi_{11}(\mathbf{x})=\prod_{i=1}^{d}f_{11}(x_{i}),\quad f_{11}(x)=\frac{\mathbb{I}\left[a\leq x\leq b\right]}{b-a}$ (19)
Unlike the previous examples, the component density $f$ and its logarithm are not continuously differentiable on $\mathbb{R}$ due to the discontinuities at the boundaries $x_{i}=a$ and $x_{i}=b$. Since this is one of the assumptions in Assumption 2.1, the asymptotic results where the optimal acceptance rate is 0.234 may not apply. This discontinuous target density case has been explored by [25], which found that for product densities with a discontinuity at the boundary, such as the unit hypercube $[0,1]^{d}$, the efficiency of the RWM is maximised with an asymptotic optimal acceptance rate of 0.1353 as $d\to\infty$. We provide simulation studies in various dimensions to investigate the applicability of this particular 0.1353 acceptance rate figure in lower dimensions on the hypercube $[0,1]^{d}$. We use 300,000 iterations instead of the typical 200,000 for this target to make the plots sufficiently smooth.
Table 5: Empirical optimal RWM acceptance rates for the continuous hypercube target.
Target $d=2$ $d=5$ $d=10$ $d=20$ $d=30$ $d=50$ $d=100$
$\pi_{11}$: Hypercube [0,1]d 0.4316 0.2767 0.2027 0.1670 0.1525 0.1439 0.1423
Compared to the previous examples (see Appendix A for their ESJD versus acceptance rate curves, which tend to be smooth and gentle at the plateau), which seem to be fairly consistently optimal with an acceptance rate of approximately 0.234, we find (refer to Table 5 and Figure 6) that the optimal acceptance rate for the hypercube target changes significantly as the dimension increases, and notice that the plateau of the curve tends to get quite sharp as the dimension increases. This is because discontinuities introduce additional complexity in the target distribution, affecting the behaviour and performance of the RWM algorithm.
Figure 6: ESJD vs. acceptance rate for the i.i.d. Hypercube $[0,1]^{d}$ target distribution $\pi_{11}$ under RWM with a Gaussian proposal in dimensions $d\in\{2,5,10,30,50,100\}$ from top-left to bottom-right. Red dotted line indicates an acceptance rate of 0.234. Purple dotted line indicates an acceptance rate of 0.135.
As the dimension $d$ increases, the algorithm’s performance converges to the theoretical 0.1353 value. In this specific case, the shrinking optimal acceptance rate might be because of the unique zero or non-zero form of the target density beyond the hypercube boundary. As discussed by [25], whether a proposed move is accepted depends completely on whether the proposed state is inside the hypercube. Subsequently, when using a Gaussian proposal, the probability that at least one proposed component will lie outside the boundary increases, leading to a higher probability of instant rejection.
The trends in Figure 6 offer practical guidance for tuning. In high dimensions (e.g., $d=100$), the ESJD curve peaks sharply near the theoretical optimum of 0.1353, and tuning to 0.234 would result in a significant loss of efficiency. However, in lower dimensions (e.g., $d=2,5,10$), the optimal acceptance rate is much closer to 0.234 than to 0.1353. The ESJD curve is also flatter, meaning the penalty for slight mistuning is less severe. This suggests a mini-max strategy for practitioners unsure of the dimensionality effects: tuning to an acceptance rate of approximately 0.234 is a robust choice across all dimensions. While it could be suboptimal for very high dimensions, it is nearly empirically optimal for low dimensions, whereas tuning to 0.1353 in these instances would be highly inefficient even though theory supports it over the 0.234 rate for this specific target distribution.
3.6 Neal’s Funnel Target Density
To probe the limits of the 0.234 heuristic, we now examine a target distribution that is canonical for its challenging geometry: the funnel distribution introduced by Neal [22]. Unlike the Rosenbrock target, which has localized curvature, Neal’s funnel possesses a global, non-linear dependency structure that presents a severe challenge to standard MCMC algorithms. In this distribution, a single variable controls the variance of all other variables, creating an extreme, state-dependent anisotropy. This structure should violate the geometric regularity conditions outlined in Assumption 2.2, particularly the target eccentricity and Hessian stability.
The $d$-dimensional Neal’s funnel target density, which we denote $\pi_{12}$, is defined for $\mathbf{x}=(x_{1},x_{2},\dots,x_{d})$. The density is given by the product of a marginal density for the first component $x_{1}$ and conditional densities for the other components:
$\pi_{12}(\mathbf{x})=\mathcal{N}(x_{1}\mid 0,3^{2})\prod_{i=2}^{d}\mathcal{N}(x_{i}\mid 0,\exp(x_{1}))$ (20)
The key feature is that the variance of each $x_{i}$ after $x_{1}$ is $\exp(x_{1})$. This creates a funnel shape in the state space: when $x_{1}$ is large and negative, the funnel is a very narrow neck where the $x_{i}$’s are tightly constrained; when $x_{1}$ is positive, the funnel opens into a wide mouth where the $x_{i}$’s can have enormous variance.
For our experiments, we follow the standard parameterization from Neal [22]. We test this target in dimensions $d\in\{5,10,20,30\}$. Due to the difficulty of sampling this target, we increased our simulation budget significantly, using $1{,}000{,}000$ iterations per simulation (across 40 seeds, 40 million total Metropolis steps), a burn-in of $50{,}000$ steps, and averaging the results over $30$ random seeds for each data point. We continue using the Gaussian proposal with variance as described in (8).
Figure 7: ESJD vs. acceptance rate for the Neal’s funnel target $\pi_{12}$ under RWM with a Gaussian proposal in dimensions $d\in\{5,10,20,30\}$ from top-left to bottom-right. Red dotted line indicates an acceptance rate of 0.234.
Similarly to the 2-dimensional Rosenbrock, we find that the ESJD increases monotonically as the acceptance rate reduces in lower dimensions. In higher dimensions ($d=20,30$), we observe a similar trend, but at the extreme ends of the curve, we observe a complete failure of the RWM algorithm to accept any reasonable number of steps as the average acceptance rate decreases towards zero, leading to sudden random drops in ESJD. Notice that the curvature of the graph gets steeper with increasing dimension: it is nearly a straight line at $d=5$ and is much more curved by $d=30$.
Overall, this behaviour indicates an extreme sensitivity to the proposal scaling in higher dimensions. The practical implication is that even minor mistuning of the proposal variance will cause the sampler’s efficiency to collapse to zero. This is a direct consequence of the target’s geometry violating the conditions of Assumption 2.2; the target has extreme eccentricity, and the Hessian is not locally stable. Consequently, the 0.234 acceptance rate is not a relevant or useful heuristic for this target. The more important conclusion, which Neal [22] himself meant to show with this example, is that for targets with strong non-linear dependencies, no single proposal scale can be efficient in both the narrow neck and the wide mouth of the funnel, leading to a sampler that inevitably fails as dimensionality grows. This experiment effectively demonstrates a hard boundary for the class of problems on which the 0.234 RWM heuristic can be successfully applied.
@@ 4.1 Parallel Tempering Method
The parallel tempering method [10], also known as the Metropolis-coupled Markov chain Monte Carlo algorithm or the replica exchange method, is a version of the Metropolis algorithm that specialises in sampling from multimodal distributions. In this case, we sample from a target distribution with density $\pi(\mathbf{x})$ on a state space $\mathcal{X}$. We define a sequence of tempered target densities $\pi^{\beta_{j}}(\mathbf{x})$ where $10\leq\beta_{n}<\beta_{n-1}<\dots<\beta_{1}<\beta_{0}=$ 1 are the inverse temperature values. We require that $\pi^{\beta_{0}}(\mathbf{x})=\pi(\mathbf{x})$ (that is, that $\beta_{0}=1$). Typically, the tempered target densities are simply powers of the original density, and we assume this for all parallel tempering experiments. Then, we run a Markov chain for each of the $n+1$ values of $\beta$, each with its own tempered target density. Eventually, the $\beta_{0}=1$ chain, also called the “cold” chain, with its stationary density $\pi^{\beta_{0}}(\mathbf{x}_{0})$ should correspond to the original target density of interest $\pi(\mathbf{x})$, and the idea is that using information from other chains with higher temperatures (lower $\beta$) will speed up the mixing in the cold chain.
In each iteration of the algorithm, we alternate between within temperature and temperature swap moves. The former indicates that each chain takes a standard RWM update within its respective $\pi^{\beta_{j}}$. The latter indicates that for two chains with $\beta_{j},\beta_{k}$ inverse temperature values that are adjacent in the sequence of $\beta$ values, switch the chain values of $\mathbf{x}_{j}$ and $\mathbf{x}_{k}$ with probability
$\min\left\{1,\dfrac{\pi^{\beta_{j}}(\mathbf{x}_{k})\pi^{\beta_{k}}(\mathbf{x}_{j})}{\pi^{\beta_{j}}(\mathbf{x}_{j})\pi^{\beta_{k}}(\mathbf{x}_{k})}\right\}.$ (21)
@@ 4.1.1 Optimal Scaling for Parallel Tempering
Since the hot chains can explore the state space more quickly, we would like to maximise how frequently we can swap values from the hottest chain to the coldest chain so that the cold chain can mix faster and escape local modes. To do this, we maximise the effective speed with which the chain values move along in the inverse temperature domain. The spacings of the inverse temperatures $\beta_{i}$’s are crucial to this efficiency. If $\beta_{j}$ and $\beta_{k}$ are too far apart, we usually reject these swaps described by (21), but if they are too close, the swaps will not improve mixing. We would like to swap just the right amount to maximise mixing in the temperature domain.
The expected squared jumping distance (ESJD) for parallel tempering thus refers to the expected squared jump in inverse temperatures. Formalising this, when we attempt to swap the chain values between the inverse temperatures $\beta$ and $\gamma:=\beta+\epsilon$ where $\beta,\epsilon>0$ and $\beta,\gamma\leq 1$, the swap is either accepted, in which case the values move a squared distance of $(\gamma-\beta)^{2}=\epsilon^{2}$, or the swap is rejected, in which case the distance moved is 0. This leads to a very natural definition for ESJD, which is
@@
There is also a 0.234 swap acceptance rate theoretical result for parallel tempering. The required assumptions for the result are the following:
Assumption Set 4.1.
(Atchadé et al. [1]) Let $\mathbf{x}$ be a random variable from the target density $\pi$.
  1. 1.
    The target density is an i.i.d. product $\pi(\mathbf{x})=\prod_{i=1}^{d}f(x_{i})$ where $f$ is a single-dimensional component density.
  2. 2.
    The tempered distributions are powers of the original density, so $\pi^{(\beta)}(\mathbf{x})=\prod_{i=1}^{d}f^{(\beta)}(x_{i})=\prod_{i=1}^{d}(f(x_{i}))^{\beta}$.
  3. 3.
    The inverse-temperature spread $\epsilon=d^{-1/2}\ell$ for some positive constant $\ell$.
Under these assumptions, choosing $\ell$ such that the inverse temperature spacing yields a swap acceptance probability of approximately 0.234 is also optimal in the sense that it maximises the ESJD of a parallel tempering algorithm.
@@ 5 Parallel Tempering Simulations
5.1 Target Distributions
5.1.1 Multivariate Gaussian Distribution
We first give an example of the standard multivariate Gaussian distribution to verify the correctness of our implementation. Note that the standard multivariate Gaussian as a target distribution satisfies the necessary conditions of Atchadé et al. [1] that are stated in Assumption 4.1.
5.1.2 Rough Carpet Distribution
In this section, we use the same target distribution $\pi_{4}$ from (12) as described in Section 3.3.1. Here, we examine the simpler case with homogeneous scaling factors $f_{d}(\mathbf{x})=\prod_{i=1}^{n}f(x_{i})$. In the individual component density $f(x)$ definition, we set $(\mu_{1},\mu_{2},\mu_{3})=(-15,0,15)$. We provide an example traceplot of the cold chain from one of the experiment simulations in Appendix A Figure 19 to demonstrate what mixing looks like in the first three components (all components are i.i.d.). Unlike the standard random-walk Metropolis which cannot escape from the central mode in this setting of the target distribution, the values in the cold chain swap between modes considerably often due to the swapping of values between chains.
5.1.3 Three Mixture Distribution
In this section, we use the same target distribution as described in Section 3.3.3 (14). We set weights $w_{1}=w_{2}=w_{3}=\frac{1}{3}$, means $\boldsymbol{\mu}_{1}=(-15,0,0,\dots,0)$, $\boldsymbol{\mu}_{2}=(0,0,\dots,0)$, $\boldsymbol{\mu}_{3}=(15,0,0,\dots,0)$ and covariances $\Sigma_{1}=\Sigma_{2}=\Sigma_{3}=I_{d}$. We provide an example traceplot of the cold chain in Appendix A Figure 20 to demonstrate what mixing looks like in the first three components.
@@ 5.2 Methodology
We conduct experiments in dimension $d\in\{10,20,30\}$ that examine the trend of ESJD with swap acceptance rate on these distributions. For each distribution, we run many simulations where each simulation is a parallel tempering algorithm with 30 different average swap acceptance rate values. We explain how we set up these algorithms to match these swap rates (through constructing a specific inverse temperature ladder) in Section 5.2.1. Each parallel tempering algorithm runs for 500,000 iterations, so every individual chain in the algorithm takes 500,000 steps. We attempt a temperature swap every 100 steps. Furthermore, we ran each algorithm instance over 20 different seeds and average the results to reduce the effects of randomness caused by a particular seed.
For each experiment, we set the tempered target distribution to simply be the original distribution raised to the power of $\beta$. We set the modes to be far away enough such that the other modes were not reached by a standard random-walk Metropolis algorithm after 100,000 iterations, but still reasonably close so that the parallel tempering method would show the cold chain values would swap between modes somewhat frequently. The within-chain proposal distribution for each chain is given by the multivariate Gaussian (8) with $\sigma^{2}=2.38^{2}/d$, which is suggested by Roberts et al. [29] to be optimal for high dimensions. We calculate the optimal swap acceptance rate by finding each simulation’s specified swap acceptance rate used to construct the iterative temperature ladder that maximises the ESJD, and take the average of the maximising acceptance rates over 20 seeds.
@@ 5.2.1 Constructing an Inverse Temperature Ladder Iteratively
The most common method to construct an inverse temperature ladder selects the inverse temperatures using a geometric series spacing. However, since we are examining the 0.234 swap acceptance rule, we need to have a way of constructing the inverse temperature spacings such that the probability of a swap between adjacent chains is approximately 0.234 (or any other value). To construct an inverse temperature ladder with a desired swap acceptance rate $s$, we use an iterative procedure adapted from [1, Section 2.2]. Starting with $\beta_{0}=1$ and a minimum value $\beta_{min}=0.01$, we iteratively add $\beta$’s to the ladder. Let $\beta_{curr}$ denote the most recent $\beta$ in the ladder. We initialize $\rho_{n}=0.5$ ($n=1$ initially) and set $\beta^{*}=\beta_{curr}(1+e^{\rho_{n}})^{-1}$. To determine if $\beta^{*}$ should be added to the ladder, we draw $N$ samples (our GPU implementation uses $N=\,$1,000,000) from the target distribution tempered by $\beta^{*}$ and $\beta_{curr}$, and calculate the average swap probability $a$. If $a$ is within $s\pm 0.0001$, we add $\beta^{*}$ to the ladder and set $\beta_{curr}=\beta^{*}$. Otherwise, we update $\rho_{n+1}=\rho_{n}+n^{-0.25}(a-s)$ and recalculate $\beta^{*}$. This process continues until $\beta^{*}\leq\beta_{min}$, at which point we add $\beta_{min}$ to the ladder and terminate.
5.3 Results
The empirical optimal swap acceptance rates for the three target distributions are summarized in Table 6, and sample ESJD curves for $d=30$ shown in Figure 8. The ESJD curves for PT are visibly more jagged than their RWM counterparts. This might be because the estimation of ESJD in this context is more susceptible to Monte Carlo error even with a large number of simulation steps, since the efficiency or acceptance rate of a swap move depends on the joint stationary distribution of two adjacent chains, and this is repeated for each temperature construction. Nonetheless, clear trends emerge from the results.
Table 6: Empirical swap acceptance rates that maximize ESJD for various PT target densities. Based on Fig 8, the jagged pattern of the ESJD curves suggests that the true optimal acceptance rates are in a neighbourhood of these values. See Section 5.2 for calculation of optimal swap acceptance rates.
Target $d=10$ $d=20$ $d=30$
$\pi_{2}$: Standard Multivariate Gaussian 0.2618 0.2685 0.2474
$\pi_{4}$: Rough Carpet 0.2660 0.2803 0.2761
$\pi_{6}$: 3-Mixture 0.2609 0.2660 0.2372
For both the standard Gaussian ($\pi_{2}$) and the non-product Three-Mixture ($\pi_{6}$) targets, the optimal swap acceptance rate is consistently in the neighbourhood of 0.234. For the Three-Mixture target, the rate converges cleanly towards the theoretical value, from 0.261 in $d=10$ to 0.237 in $d=30$. The Gaussian target shows a similar pattern, with an optimal rate of 0.247 at d=30.
The standard Gaussian result provides a successful validation of our experimental setup against the theory of Atchadé et al. [1], as this target satisfies the i.i.d. product form assumption. More importantly, the result for the Three-Mixture target suggests that the 0.234 heuristic for PT swaps is robust to violations of the i.i.d. product assumption, much like its RWM counterpart. Although the target is a sum of densities, its underlying geometric regularity (a shell composed of three isotropic modes) seems sufficient for the optimal scaling 0.234 result to hold. For a practitioner, this suggests that constructing a temperature ladder with a swap acceptance probability of 0.234 is a viable and likely near-optimal strategy, even for complex, non-product multimodal targets, provided they are not pathologically structured.
(a) $\pi_{2}$: Gaussian
(b) $\pi_{4}$: Rough Carpet
(c) $\pi_{6}$: 3-Mixture